1. A car with tires of diameter 787 mm travels on the highway at 76 mph (1 mph = 0.44704 m/s). How long does it take a pebble stuck on the tread of the tire on its radial edge to revolve around the tire?

Report the value in milliseconds rounded to the nearest whole number.
1
Expert's answer
2020-03-25T11:22:24-0400
"v=\frac{2\pi r}{T}\to T=\frac{2\pi r}{v}"

"T=\frac{2\pi (0.787)}{(76)(0.44704)}=0.146\ s=146\ ms"